A valószínűségszámítás I. gyakorlat április 2. dolgozatának eredményei.

dolgozat példák: 1. 2 . 3. 4. 5. 6. gyakorlati

jegy

diákok:

Csaba Anikó  −  +  ±  +  +  −  4
Dorogi Edina  ±  +  ±  +  +  +  3
Jámbor Tamás  −  +  +  +  +  +  4
Medve Diana  −  +  −  −  −  +  —
Mészáros Fruzsina  −  +  +  +  +  +  5
Ölbei Kinga  −  ±  +  ±  +  +  2
Suta Judit  −  +  +  +  +  +  4
Tanka Irén  −  ±  +  ±  +  +  2
Tóth Boglárka  −  +  +  −  +  −  3